# On the gravitational stratification of multi-fluid-multi-species plasma

### 17 Mar 2026 v2

@@ Abstract
Results. A gravitational stratification constructed using the present numerical integration routine can be in hydrostatic equilibrium with any given ionization fractions of multi-species plasmas. Meanwhile, without any dynamic driving force, fluid decoupling appears in the transition region of the constructed stratification, while the total velocity of all fluids remains zero.
@@ 1 Introduction
where total pressure includes thermal pressure of gas plasma and magnetic pressure, i.e., $P=p^{\text{g(as)}}+p^{\text{m(agnetic)}}$, $\varrho$ is plasma density, $g$ is gravitational acceleration, and $z$ is height. This balance between the total pressure gradient and gravity is assumed to hold in the atmospheric model along the $z$-direction. Note that, in the solar atmosphere, the mean free path of particles is much smaller than the characteristic length scales of interest. Therefore, the plasma can be assumed to be in Maxwellian equilibrium, and a fluid description, e.g., magnetohydrodynamics (MHD), is appropriate. Consequently, the stationary momentum equation reduces to the hydrostatic balance as written in Eq. (1). Models with this assumption have been extensively used to investigate, for example, MHD waves in the solar atmosphere, being able to provide important insights (Khomenko & Collados, 2015; Van Doorsselaere et al., 2020).
@@
in which the thermal pressure $p_{i}^{\text{g}}$ of the $i$-th fluid includes the thermal pressure of the free electrons produced when the $i$-th fluid undergoes ionization. For example, in various pure-hydrogen ion-neutral two-fluid models (Alvarez Laguna et al., 2016; Popescu Braileanu et al., 2019; Wójcik et al., 2020; Popescu Braileanu & Keppens, 2023), the equations of state of neutral and ionized hydrogen are, respectively,
@@
where $N$ and $T$ denote the number densities and temperatures, with subscripts i, n, and e indicating ionized hydrogen, neutral hydrogen, and free electrons, respectively; $k_{\text{B}}$ is the Boltzmann constant. These two independent equations of state lead to different scale heights of neutral and ionized hydrogen. Moreover, in these models, the description of free electrons is simplified, ignoring the thermal force, which is feasible for modeling the low solar atmosphere which can be assumed to be nearly isothermal (Hansteen et al., 1997). In addition, Eq. (4) assumes that the ion temperature and electron temperature are the same. This assumption simplifies the numerical model but is not a necessary constraint for the solar atmosphere.
@@
Directly using Eq. (2) to calculate a multi-fluid gravitational stratification causes practical issues. As the scale heights are different for different fluids, the resulting abundances at certain heights may differ by orders of magnitude from atmospheric models that holistically combine the contributions of necessary physical mechanisms (Vernazza et al., 1981; Fontenla et al., 1993; Avrett & Loeser, 2008). Such differences then change plasma properties and thus the results from numerical models (Zhang et al., 2021). Moreover, what Eq. (1) or (2) describes is a boundary value problem that depends on the values of $\varrho_{i}(z_{0})$, in which $z_{0}$ is the reference height, and temperature only plays a role through the equations of state. Thus the resulting stratified plasma is largely not in ionization equilibrium, which directly relates to the local temperature, and static atmospheric models given by Eq. (2) cannot be used to start simulations with ionization and recombination, which will drastically change the energy balance. Using Eq. (2) also implies the assumption that collisional interactions between different fluids do not affect the static equilibrium. Intuitively, however, one may still expect collisional interactions between fluids, although such interactions may not always be sufficient to fully couple all fluids (Hansteen et al., 1997). Therefore, it is necessary to revise the assumption of independent hydrostatic equilibria for different fluids.
@@
Obtaining an initial atmospheric model that is not only in hydrostatic equilibrium but also in ionization equilibrium, thus far, is not trivial, and only a few two-fluid low solar atmospheric models (Maneva et al., 2017; Zhang et al., 2021; Niedziela et al., 2024; Kraśkiewicz & Murawski, 2025) have considered both equilibria. Indeed, it is possible to run a long(er) physical-time simulation to reach a steady state (Brchnelova et al., 2023), in which the initial imbalances eventually disappear. However, the numerical solution depends on the chosen initial conditions, and it is difficult to constrain the solution to obtain a specific atmospheric composition. In addition, this strategy can also be computationally intensive. Therefore, to obtain an initial static stratification for a multi-fluid-multi-species (MFMS) MHD model that may include various heavier elements (Wargnier et al., 2023; Martínez-Sykora et al., 2023) in addition to hydrogen, an alternative solution is needed.
@@
To construct an MFMS hydrostatic equilibrium atmospheric model, we assume a static (or steady) configuration in the absence of external disturbances or force imbalances, such that the characteristic timescale of the equilibrium state is infinitely larger than the collisional timescales. Therefore, we may assume that all the $n$ fluids are coupled by collisional interactions (or other mixing processes) [^1] This assumption is, however, not valid higher up in the solar atmosphere and the solar wind, where the plasma gradually becomes collisionless and different fluids/species become decoupled (Hansteen et al., 1994; Hansteen & Leer, 1995). The present MFMS model is expected to be used for the low solar atmosphere., approximately acting as a single fluid (Khomenko & Collados, 2012; Martínez-Sykora et al., 2012) in the static state. Then, the gravitational stratification can be calculated straightforwardly. Zhang et al. (2021) used this idea to provide initial equilibria for pure-hydrogen two-fluid plasmas under an isothermal condition. The idea can be extended to MFMS modeling by using a simple numerical integration routine, which is the focus of this work. Below, we begin by explaining this simple idea using an isothermal example similar to those of Zhang et al. (2021), and then more realistic temperature stratification and ionization are taken into account. We also address the essential physical and numerical consequences of using our model. Additional details are introduced in the Appendix.
2 An isothermal example
To explain the idea of this work and its necessity, we extend the solution given by Zhang et al. (2021). We further include neutral and ionized helium (He and He+), in addition to neutral and ionized hydrogen (H and H+), so the model has two species and four fluids (not counting free electrons).
We assume an upper chromospheric condition adopted from a 2.5D radiative MHD (rMHD) model (Martínez-Sykora et al., 2020a) using the Bifrost code (Gudiksen et al., 2011). The condition was found in regions where the ponderomotive force may produce chemical fractionation, associated with spicules or low-lying loops (Martínez-Sykora et al., 2023). Specifically, the temperature considered is $1.6\times 10^{4}$ K, and the number densities of H, H+, He, and He+ were found to be respectively 109.8, 1010.7, 109.7, and 108.3 per cm3. These number densities are a result of solving for nonequilibrium ionization (NEQ), i.e., solving the full time-dependent rate equations for hydrogen and helium. When, instead, the rMHD model assumes statistical equilibrium (SE) ionization, i.e., solving the rate equations without any $\partial N_{i}/\partial t$ term, where $N_{i}$ is the number density of the $i$-th fluid, the composition typically includes significantly more neutral fluid.
@@ 2 An isothermal example
Based on the given condition, we calculate and discuss two different types of stratification:
@@
As we assume that in the cHE stratification all fluids are coupled by collisional interactions, at each height all fluids can be described by their bulk motion. We define a mean atomic mass [^2] Hydrogen molecules are largely dissociated at the given temperature. for considering the bulk motion, which reads
@@
where $m_{i}$ is the atomic mass of the $i$-th fluid. We should particularly emphasize that, free electrons should be considered (thus the nn+1-th fluid) when calculating the mean atomic mass, while the electron mass is assumed to be zero. Note that we do include separate equations to describe the motion of electrons, as further explained in Section 4. While assuming the ideal gas law, we have the equivalent equation of state
@@
assuming that the temperature is the same for all fluids. Finally, Eq. (1) can be rewritten for the MFMS model, giving us
$\frac{\text{d}P}{\text{d}z}=-N_{\text{T}}\overline{m}_{\text{T}}g=-\frac{p_{\text{T}}^{\text{g}}\overline{m}_{\text{T}}g}{k_{\text{B}}T}.$ (8)
For the non-magnetic and isothermal ($T\equiv 1.6\times 10^{4}$ K) case here, an analytical solution
@@
Note that the mean atomic mass is used for calculating the plasma gas pressure. When calculating the (mass) density of a specific fluid, we simply use its number density to multiply the real atomic mass of the fluid. In the present scenario, the ionization fractions do not change over height in the cHE stratification, and thus the (partial) number densities of all fluids can be easily calculated after the total gas pressure has been integrated.
In Figure 1, we compare the pHE and cHE results given within a 1 Mm interval [^3] This interval is not suggested to be a realistic layer of the solar atmosphere., using the aforementioned NEQ condition to provide the plasma quantities at $z=0$. The pHE solution results in exponentially changing ionization fractions, as each fluid has its own scale height, in contrast to the cHE solution. Moreover, in the cHE solution, each fluid is not in pure hydrostatic balance. Instead, gravity is balanced by the combination of the pressure gradient and the collisional forces, assuming that collisions are sufficiently strong. Therefore, in a cHE stratification, the velocity of each fluid is not zero, but their total velocity ($w_{\text{T}}$) along the $z$ direction
@@
is expected to be approximately zero, within the integration error of Eq. (8). A similar form of Eq. (10) is introduced by Khomenko et al. (2014) as the velocity of the center of mass in the one-fluid model, which also explains the static assumption we use before any dynamic driving force is imposed. However, we do not include electrons when defining the total velocity, as the electron velocity is defined as a function of velocities of ions and the total current, i.e., Eq. (13).
@@ 3 General scenarios
3.1 A simple numerical integration routine
In the previous section, we used a simple example to analytically construct a coupled MFMS stratification that is in balance between gravity, pressure gradients, and collisional forces between multiple fluids/species. In more general scenarios, in which temperature is not constant, ionization fractions should vary even if SE is assumed. Therefore, a numerical routine is needed to integrate Eq. (8).
Below, we assume that the ionization fractions of all species are already calculated over the entire 1D domain based on any given temperature distribution, and are not constant. The ionization fractions can be given by assuming the SE condition under realistic solar atmospheric temperature distributions, or provided by self-consistent rMHD simulations with, for instance, NEQ ionization. We then discretize Eq. (8) and obtain a numerical integration formula
$P_{j+1}=P_{j}-\frac{p_{\text{T},j}^{\text{g}}\overline{m}_{\text{T},j}g_{j}}{k_{\text{B}}T_{j}}\Delta z_{j},\quad j=0,...,M,$ (11)
where $M$ is the number of intervals, with $\Delta z_{j}=z_{j+1}-z_{j}$, $\overline{m}_{\text{T},j}$ is calculated using Eq. (5), and $T_{j}$ is given and not changed by the integration routine. $\overline{m}_{\text{T},j}$ and $T_{j}$ are also considered to be constant in each integration step. More specifically, at discrete point $j$, the local ionization fraction is used to calculate local partial number densities, which are then used in Eq. (5). Except at $j=0$, where the boundary solution is imposed, the local total number density is calculated using (I) the total gas pressure provided by the integration step j1j-1, and (II) the equivalent equation of state, Eq. (6). Note that the contribution of magnetic pressure is discussed in Appendix A.
In practice, $M$ is preferably larger than the number of grid points used in numerical simulations, to reduce the error in the obtained stratification. This is beneficial as the present integration routine is effectively a first-order left Riemann sum. The present solution is intuitively straightforward, as each integration step is an isothermal scenario introduced in the last section. However, the assumptions we use also mean that neither the hydrostatic equilibrium nor the ionization equilibrium is strictly maintained between any pair of neighboring numerical grid points due to numerical error. Moreover, while collisional interactions are needed to balance gravity and pressure gradient, there must be drift velocities between fluids. Only the total velocity of all fluids is expected to be approximately zero, as has been mentioned above.
@@ 3.2 Revisiting the SE assumption
We calculate an SE stratification by first extracting the temperature distribution from the widely used Model C7 (Avrett & Loeser, 2008). The total number density at the bottom of the photosphere ($z=0$ Mm) is approximated using the hydrogen number density given by Model C7, as ion and electron number densities are still small on the solar surface. In addition to the four fluids included in Section 2, we further include [^4] There is effectively no limitation on the number of species that Ebysus can include, as long as relevant information (e.g. collisional cross sections) are available. Different species are included in this work simply to show the flexibility of the present approach. Ne, Ne+, Fe, and Fe+. Then, SE ionization under the given temperature distribution is calculated to provide the number densities of these eight fluids, using the ionization and recombination rates of Voronov (1997) and McWhirter (1965), with photospheric abundances of Asplund et al. (2021) provided through the CHIANTI atomic database, Version 11.0 (Dufresne et al., 2024). Note that we calculate only the cHE distributions here, because pHE would lead to significant deviations in the resulting ionization fractions, as explained in the previous section. The cHE stratification is shown in Figure 2.
We find that the total ion density below the transition region (TR) is around 10 orders of magnitude lower than the total neutral density, which relates to several factors. Firstly, the ionization state of hydrogen and helium is computed in SE, including collisional ionization balanced by radiative and three-body recombinations (for heavier ions, dielectronic recombinations would also be included), using rates from the CHIANTI package, while leaving out photoionizations. In addition, as mentioned, SE may underestimate ion fractions due to slow ionization and recombination rates for some transitions. Thus, in the upper chromosphere, NEQ ionization is required, while in the photosphere, solving the Saha equation (Saha, 1920) including all species is needed, as metals with their low FIP (first ionization potential) are the main donors of electrons. A more realistic calculation of SE ionization was presented by Gómez-Míguez et al. (2024), who used the MULTI code (Carlsson, 1986) to solve the radiative transfer equation together with the rate equations, thus including photoionization correctly. However, how the ionization fractions are calculated does not affect the numerical integration. Figure 2 shows that, over the whole domain, the ionization fractions are in SE. Therefore, the present cHE stratification is in both hydrostatic equilibrium and SE.
@@
The previous stratification is calculated using ionization fractions based on the ionization and recombination rates in SE. However, NEQ ionization fractions can also be used to calculate the number densities needed for the numerical integration. Below we prove the flexibility of the present numerical integration routine, which is applicable as long as temperature and ionization fraction distributions are provided, irrespective of how they are calculated.
We adopt a NEQ simulation snapshot from the 2.5D rMHD model (Martínez-Sykora et al., 2020a), as shown in Figure 3. In this model, we observe $\sim$50 Mm long loops, which are associated with the underlying magnetic field configuration, which has two main opposite polarities. Correspondingly, there are cold and dense features with chromospheric temperature along the loops, resembling spicules and jets (Martínez-Sykora et al., 2017). More details of the rMHD simulation are not further discussed here. We calculate the horizontally-averaged density and temperature, and then at first use the method introduced in Section 3.2 to obtain the SE ionization fractions. Note that the result here includes two-times ionized helium (He2+) but not iron or neon, resulting in a two-species-five-fluid mixture. Subsequently, for comparison, we also extract the averaged NEQ ionization fractions from the 2.5D snapshot, replacing the SE ionization fractions, to perform the same integration routine. We obtain two stratifications shown in Figure 4, while omitting the contribution of magnetic pressure gradient. The inclusion of magnetic pressure is briefly discussed in Appendix A. Note that both stratifications are calculated using the same temperature distribution.
@@
Due to the horizontal averaging, we obtain a relatively flat TR from the rMHD model. Still, the aim is to show that a cHE stratification can be obtained to include the effects of NEQ ionization. Specifically, the NEQ stratification has higher ion populations in the chromosphere and photosphere, resulting from the long time-scales of the ionization and recombination processes, as well as their interplay with chromospheric dynamics, e.g., shocks (Carlsson & Stein, 1992, 2002; Leenaarts et al., 2007; Golding et al., 2014; Wedemeyer-Böhm & Carlsson, 2011). Therefore, the present numerical integration routine provides flexibility for specifically studying atmospheric dynamics while considering the consequence of NEQ ionization. In contrast, a pHE stratification cannot recover such effects, as shown in Figure 1.

### 30 Apr 2026 v3

@@ Abstract
Results. A gravitational stratification constructed using the present numerical integration routine can be in hydrostatic equilibrium with any given ionization fractions of multi-species plasmas. Meanwhile, without any dynamic driving force, fluid decoupling appears, particularly in the transition region of the constructed stratification, while the total velocity of all fluids remains zero.
@@ 1 Introduction
where total pressure includes thermal pressure of gas plasma and magnetic pressure, i.e., $P=p^{\text{g(as)}}+p^{\text{m(agnetic)}}$, $\varrho$ is plasma density, $g$ is gravitational acceleration, and $z$ is height. This balance between the total pressure gradient and gravity is assumed to hold in the atmospheric model along the $z$-direction. Note that, in the solar atmosphere, the mean free path of particles is much smaller than the characteristic length scales of interest. Therefore, the plasma can be assumed to be in Maxwellian equilibrium, and a fluid description, e.g., magnetohydrodynamics (MHD), is appropriate. Consequently, the stationary momentum equation reduces to the hydrostatic balance as written in Eq. (1). Models with this assumption have been extensively used to investigate, for example, MHD waves in the solar atmosphere (Khomenko and Collados, 2015; Van Doorsselaere et al., 2020).
@@
in which the thermal pressure $p_{i}^{\text{g}}$ of the $i$-th fluid includes the thermal pressure of the free electrons produced when the $i$-th fluid underwent ionization. For example, in various pure-hydrogen ion-neutral two-fluid models (Alvarez Laguna et al., 2016; Popescu Braileanu et al., 2019; Wójcik et al., 2020; Popescu Braileanu and Keppens, 2023), the equations of state of neutral and ionized hydrogen are, respectively,
@@
where $N$ and $T$ denote the number densities and temperatures, with subscripts i, n, and e indicating ionized hydrogen, neutral hydrogen, and free electrons, respectively; $k_{\text{B}}$ is the Boltzmann constant. These two independent equations of state lead to different and independent scale heights of neutral and ionized hydrogen. Moreover, in these models, the description of free electrons is simplified, ignoring the thermal force, which is feasible for modeling the low solar atmosphere which can be assumed to be nearly isothermal (Hansteen et al., 1997). In addition, Eq. (4) assumes that the ion temperature and electron temperature are the same. This assumption simplifies the numerical model but is not a necessary constraint for the solar atmosphere.
@@
Directly using Eq. (2) to calculate a multi-fluid gravitational stratification causes practical issues. As the scale heights are independent for different fluids, the resulting abundances at certain heights may differ by orders of magnitude from atmospheric models that holistically combine the contributions of necessary physical mechanisms (Vernazza et al., 1981; Fontenla et al., 1993; Avrett and Loeser, 2008). Such differences then change plasma properties and thus the results from numerical models (Zhang et al., 2021). Moreover, what Eq. (1) or (2) describes is a boundary value problem that depends on the values of $\varrho_{i}(z_{0})$, in which $z_{0}$ is the reference height, and temperature only plays a role through the equations of state. Thus, the resulting stratified plasma is largely not in ionization equilibrium, which directly relates to the local temperature, and static atmospheric models given by Eq. (2) cannot be used to start simulations with ionization and recombination, which will drastically change the energy balance. Using Eq. (2) also implies the assumption that collisional interactions between different fluids do not affect the static equilibrium. Intuitively, however, one may still expect collisional interactions between fluids, although such interactions may not always be sufficient to fully couple all fluids (Hansteen et al., 1997). Physically, collisional processes are also important in ionization equilibrium (House, 1964). Therefore, it is necessary to revise the assumption of independent hydrostatic equilibria for different fluids.
@@
Obtaining an initial atmospheric model that is not only in hydrostatic equilibrium but also in ionization equilibrium, thus far, is not trivial, and only a few two-fluid low solar atmospheric models (Maneva et al., 2017; Zhang et al., 2021; Niedziela et al., 2024; Kraśkiewicz and Murawski, 2025) have considered both equilibria. Indeed, it is possible to run a long(er) physical-time simulation to reach a steady state (Brchnelova et al., 2023), in which the initial imbalances eventually disappear. However, the numerical solution would depend on the chosen initial conditions, and it is difficult to constrain the solution to obtain a specific atmospheric composition. In addition, this strategy can also be computationally intensive. Therefore, to obtain an initial static stratification for a multi-fluid-multi-species (MFMS) model that may include various heavier elements (Wargnier et al., 2023; Martínez-Sykora et al., 2023) in addition to hydrogen, an alternative solution is needed.
@@
To construct an MFMS hydrostatic equilibrium atmospheric model, we assume a static (or steady) configuration in the absence of external disturbances or force imbalances, such that the characteristic timescale of the equilibrium state is infinitely larger than the collisional timescales. Therefore, we may assume that all the $n$ fluids are coupled by collisional interactions (or other mixing processes) [^1] This assumption is, however, not valid higher up in the solar atmosphere and the solar wind, where the plasma gradually becomes collisionless and different fluids/species become decoupled (Hansteen et al., 1994; Hansteen and Leer, 1995)., approximately acting as a single fluid (Khomenko and Collados, 2012; Martínez-Sykora et al., 2012) in the static state. Then, the gravitational stratification can be calculated straightforwardly. Zhang et al. (2021) used this idea to provide initial equilibria for pure-hydrogen two-fluid plasmas under an isothermal condition. The idea can be extended to MFMS modeling by using a simple numerical integration routine, which is the focus of this work. Below, we begin by explaining this simple idea using an isothermal example similar to those of Zhang et al. (2021), and then more realistic temperature stratification and ionization are taken into account. We also address the essential physical and numerical consequences of using our model. Additional details are introduced in the Appendix.
2 The basics
2.1 A coupled approach
@@
Considering multi-fluid plasmas, we calculate and discuss two different types of stratification:
@@
As we assume that in the cHE stratification all fluids are coupled by collisional interactions, at each height all fluids can be described by their bulk motion. We thus define a mean atomic mass, which reads
@@
where $m_{i}$ and $N_{i}$ are respectively the atomic mass and the number density of the $i$-th fluid. We should particularly emphasize that, free electrons should be considered (thus the nn+1-th fluid) when calculating the mean atomic mass, without breaking charge neutrality, while the electron mass is assumed to be zero. While assuming the ideal gas law, we have the equivalent equation of state
@@
by assuming that the temperature is the same for all fluids. Finally, Eq. (1) can be rewritten for the MFMS model, giving us
$\frac{\text{d}p_{\text{T}}^{\text{g}}}{\text{d}z}=-N_{\text{T}}\overline{m}_{\text{T}}g=-\frac{p_{\text{T}}^{\text{g}}\overline{m}_{\text{T}}g}{k_{\text{B}}T},$ (8)
without accounting for the contribution of magnetic pressure. Magnetic pressure (gradient) can be included straightforwardly, as described in the first section and Appendix A.
Note that the formulae above are valid for the cHE stratification in general scenarios.
2.2 An isothermal example
We extend the solution given by Zhang et al. (2021), by further including neutral and ionized helium (He and He+), in addition to neutral and ionized hydrogen (H and H+). Therefore, this example has two species and four fluids (not counting free electrons).
More specifically, we assume an upper chromospheric condition adopted from a 2.5D radiative MHD (rMHD) model (Martínez-Sykora et al., 2020a) using the Bifrost code (Gudiksen et al., 2011). The condition was found in regions where the ponderomotive force may produce chemical fractionation, associated with spicules or low-lying loops (Martínez-Sykora et al., 2023). Specifically, the temperature considered is $1.6\times 10^{4}$ K, and the number densities of H, H+, He, and He+ were respectively 109.8, 1010.7, 109.7, and 108.3 per cm3. These number densities are a result of solving for nonequilibrium ionization (NEQ), i.e., solving the full time-dependent rate equations for hydrogen and helium. When, instead, the rMHD model assumes statistical equilibrium (SE) ionization (House, 1964; McWhirter, 1965; Voronov, 1997), i.e., solving the rate equations without any $\partial N_{i}/\partial t$ term, the composition typically includes significantly more neutral fluids.
Then, for a non-magnetic and isothermal ($T\equiv 1.6\times 10^{4}$ K) case, an analytical solution
@@ 2.2 An isothermal example
Note that the mean atomic mass is used for balancing the plasma gas pressure gradient in Eq. (8). When calculating the (mass) density of a specific fluid, we simply use its number density to multiply the real atomic mass of the fluid. In the present scenario, the ionization fractions do not change over height in the cHE stratification, and thus the (partial) number densities of all fluids can be easily calculated after the total gas pressure has been integrated.
In Figure 1, we compare the pHE and cHE results given within a 1 Mm interval [^2] This interval is not suggested to be a realistic layer of the solar atmosphere., using the aforementioned NEQ condition to provide the plasma quantities at $z=0$. The pHE solution results in exponentially changing ionization fractions, simply because of gravitational stratification. In the cHE solution, each fluid is not in pure hydrostatic balance. Instead, gravity is balanced by the combination of the pressure gradient and the collisional forces, assuming that collisions are sufficiently strong.
In a cHE stratification, the velocity of each fluid may not be zero, but their total velocity ($w_{\text{T}}$) along the $z$ direction
@@
is expected to be within the integration error of Eq. (8). Specifically, when the temperature is constant, an analytical solution can be used and thus there is no error. In the following we discuss general cases. A similar form of Eq. (10) is introduced by Khomenko,E. et al. (2014) as the velocity of the center of mass in the one-fluid model, which also explains the static assumption we use before any dynamic driving force is imposed. However, we do not include electrons when defining the total velocity, as the electron velocity is defined as a function of velocities of ions and the total current, i.e., Eq. (13).
@@ 3 General scenarios
3.1 A numerical integration routine
In the previous section, we constructed a coupled MFMS stratification that is in balance between gravity, pressure gradients, and collisional forces between multiple fluids/species. In more general scenarios, in which temperature is not constant, ionization fractions should vary even if SE is assumed. Therefore, a numerical routine is needed to integrate Eq. (8).
Below, we assume that the temperature distribution is provided. The ionization fractions can be calculated by assuming the SE condition under realistic solar atmospheric temperature distributions, or provided by self-consistent rMHD simulations with, for instance, NEQ ionization. Then, firstly, we discretize Eq. (8) in the 1D domain and obtain a numerical integration formula starting from the lower boundary of the domain
$p_{\text{T},j+1}^{\text{g}}=p_{\text{T},j}^{\text{g}}-\frac{p_{\text{T},j}^{\text{g}}\overline{m}_{\text{T},j}g_{j}}{k_{\text{B}}T_{j}}\Delta z_{j},\quad j=0,...,M,$ (11)
where $M$ is the number of intervals, with $\Delta z_{j}=z_{j+1}-z_{j}$, $\overline{m}_{\text{T},j}$ is calculated using Eq. (5), for which the partial number densities are needed, and $T_{j}$ is given and not changed by the integration routine. $\overline{m}_{\text{T},j}$ and $T_{j}$ are considered to be constant within each integration interval.
At discrete point $j$, if the local ionization fraction is not provided in advance, it can be calculated using the given local temperature. Except at $j=0$, where the boundary condition is imposed, the local total gas pressure $p_{\text{T},j}^{\text{g}}$ is provided by integration step j1j-1. The local total number density is calculated using the total gas pressure and the equivalent equation of state [^3] When the Saha-Boltzmann equation is used, the local electron pressure can also be calculated by the j1j-1 step, to provide the local electron number density., Eq. (6). The ionization fraction and the total number density together provide local partial number densities, with which $\overline{m}_{\text{T},j}$ can be calculated. With $p_{\text{T},j}^{\text{g}}$ and $\overline{m}_{\text{T},j}$ provided, $p_{\text{T},j+1}^{\text{g}}$ can be calculated.
In practice, $M$ is preferably larger than the number of grid points used in numerical simulations. This is because the present integration routine is effectively a first-order left Riemann sum, and thus the integration intervals need to be sufficiently small to reduce numerical errors. The present solution is intuitively straightforward, as each integration step is an isothermal scenario introduced in the last section. However, the assumptions we use also mean that neither the hydrostatic equilibrium nor the ionization equilibrium is strictly maintained between any pair of neighboring grid points due to numerical errors. Moreover, while collisional interactions are needed to balance gravity and pressure gradient, there must be drift velocities between fluids. Only the total velocity of all fluids is expected to be approximately zero, as has been mentioned above.
@@ 3.2 Revisiting the SE assumption
We calculate an SE stratification by first extracting the temperature distribution from the widely used Model C7 (Avrett and Loeser, 2008). The total number density at the bottom of the photosphere ($z=0$ Mm) is approximated using the hydrogen number density given by Model C7, as ion and electron number densities are still small on the solar surface. In addition to the four fluids included in Section 2, we further include [^4] There is effectively no limitation on the number of species that Ebysus can include, as long as relevant information (e.g. collisional cross sections) are available. Different species are included in this work simply to show the flexibility of the present approach. Ne, Ne+, Fe, and Fe+. Then, SE ionization under the given temperature distribution is calculated to provide the number densities of these eight fluids, using the ionization and recombination rates of Voronov (1997) and McWhirter (1965), with photospheric abundances of Asplund et al. (2021) provided through the CHIANTI atomic database, Version 11.0 (Dufresne et al., 2024). Note that we calculate only the cHE distributions here, because pHE would lead to significant deviations in the resulting ionization fractions, as explained in the previous section. The cHE stratification is shown in Figure 2.
We find that the total ion density below the transition region (TR) is around 10 orders of magnitude lower than the total neutral density, which relates to several factors. Firstly, the ionization state of hydrogen and helium is computed in SE, including collisional ionization balanced by radiative and three-body recombinations (for heavier ions, dielectronic recombinations would also be included), using rates from the CHIANTI package, while leaving out photoionizations. In addition, as mentioned, SE may underestimate ion fractions due to slow ionization and recombination rates for some transitions. Thus, in the upper chromosphere, NEQ ionization is required, while in the photosphere, solving the Saha equation (Saha, 1920; Rutten, 2003) including all species is needed, as metals with their low FIP (first ionization potential) are the main donors of electrons. A more realistic calculation of SE ionization was presented by Gómez-Míguez et al. (2024), who used the MULTI code (Carlsson, 1986) to solve the radiative transfer equation together with the rate equations, thus including photoionization correctly. However, how the ionization fractions are calculated does not affect the numerical integration. Figure 2 shows that, over the whole domain, the ionization fractions are in SE. Therefore, the present cHE stratification is in both hydrostatic equilibrium and SE.
@@
The previous stratification is calculated using ionization fractions based on the ionization and recombination rates in SE. However, NEQ ionization fractions can also be used to calculate the number densities needed for the numerical integration. Below we prove the flexibility of the present numerical integration routine, which is applicable as long as temperature and ionization fraction distributions are provided.
We adopt a NEQ simulation snapshot from the 2.5D rMHD model (Martínez-Sykora et al., 2020a), as shown in Figure 3. In this model, we observe $\sim$50 Mm long loops, which are associated with the underlying magnetic field configuration, which has two main opposite polarities. Correspondingly, there are cold and dense features with chromospheric temperature along the loops, resembling spicules and jets (Martínez-Sykora et al., 2017). More details of the rMHD simulation are not further discussed here. We calculate the horizontally-averaged density and temperature, and then at first use the method introduced in Section 3.2 to obtain the SE ionization fractions. Note that the result here includes two-times ionized helium (He2+) but not iron or neon, resulting in a two-species-five-fluid mixture. Subsequently, for comparison, we also extract the averaged NEQ ionization fractions from the 2.5D snapshot, replacing the SE ionization fractions, to perform the same integration routine. We obtain two stratifications shown in Figure 4, while omitting the contribution of magnetic pressure. The inclusion of magnetic pressure is discussed in Appendix A. Note that both stratifications are calculated using the same temperature distribution.
@@
Due to the horizontal averaging, we obtain a relatively flat TR from the rMHD model. Still, the aim is to show that a cHE stratification can be obtained to include the effects of NEQ ionization. Specifically, the NEQ stratification has higher ion populations in the chromosphere and photosphere, resulting from the long time-scales of the ionization and recombination processes, as well as their interplay with chromospheric dynamics (Carlsson and Stein, 1992, 2002; Leenaarts et al., 2007; Wedemeyer-Böhm and Carlsson, 2011; Golding et al., 2014). Therefore, the present numerical integration routine provides flexibility for specifically studying atmospheric dynamics while considering the consequence of NEQ ionization. In contrast, a pHE stratification cannot recover such effects, as shown in Figure 1.